# AICO: Feature Significance Tests for Supervised Learning

### v2

@@ 2.1 Baseline and comparison features
A few basic principles guide the selection of reference points and the construction of the set $\underline{I}\subseteq I_{1}$. We wish to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, to avoid feature effects $\Delta_{i}$ that are biased low which result in significance tests with low power.555? discuss these issues in the context of image classification. Similarly, we want to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$. Finally, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space. It is important to select an appropriate reference point and rule used to aggregate across $\underline{I}$ (see Example 2.1). Alternatively, we might select an appropriate baseline value from the set $\{X_{k}:k\in I_{1}\}$, namely an element that is near the raw off-manifold value emerging from the baseline selection procedure.
For clarity, the discussion above has focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$. The approach naturally extends to testing multiple features, i.e. subsets of $[d]$, which facilitates the exploration of higher-order effects including feature interactions. We also allow for the selection and aggregation of alternative baselines for a given sample. A general tool is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
2.2 Panel data
Panel data settings, where a collection of units are observed over some time period, are prevalent in many domains. While our approach assumes independent samples, a simple extension—grouping—allow us to treat some dependent settings like panel data. The idea is to aggregate feature effects within each unit of dependence, so that dependence is contained within groups with inter-group dependencies being mitigated.
Concretely, consider panel data $\{Z_{i,t}\}$, where $i=1,\dots,m$ indexes units and $t=1,\dots,T$ indexes time periods. We take as given a unit-level random split of the index set $m]=\{1,\dots,m\}$ into disjoint training and testing subsets $I_{1}$ and $I_{2}$, respectively, such that each unit, along with its entire time series, is assigned exclusively to either $I_{1}$ or $I_{2}$. We treat the test set $I_{2}$ as a collection of $|I_{2}|T$ samples and compute the per-unit-time feature effects $\Delta_{i,t}$ for each $(i,t)\in I_{2}$. We then aggregate these effects within each unit: $\Delta_{i}=\frac{1}{T}\sum_{t=1}^{T}\Delta_{i,t}$.
For purposes of computing baseline and comparison features in this group setting, we also treat the training data $I_{1}$ at a unit level. For each $(i,t)\in I_{1}$ and feature $j\in[d]$, the per-unit-time samples are aggregated to obtain grouped feature samples $X_{i}^{j}=a_{i}^{j}(X^{j}_{i,1},\dots,X^{j}_{i,T})$, where $a_{i}^{j}$ is an aggregation function for the time series associated with unit $i$ (e.g. the mean, median, mode).
The grouping approach naturally extends to unbalanced panel data, where subjects may be observed at different time points or for different durations. In such cases, the average for each subject is computed only over the observed periods.
More broadly, the grouping approach can be adapted to other dependence structures beyond temporal ones. For instance, for spatially dependent data, one could group feature effects by region or neighborhood, under the assumption that neighborhoods are (approximately) independent. Similarly, in settings with repeated measurements, grouping by the underlying entity or context is effective.
@@ 5.1 Regression
We run 10 independent training/testing experiments to assess the behavior of the test. The second and third columns of Table 1 report the results for confidence levels $\alpha=0.01$ and 0.050.05.888Limited numerical precision can lead to ties, i.e. samples $\Delta_{i}$ whose true values are distinct are being rounded to the same value. If the value of a tie is strictly positive, we count each sample in the test statistic $n_{+}$. Another numerical artifact is samples $\Delta_{i}$ with true, non-zero value being rounded to 0. We treat these samples as if they indeed had non-positive values not contributing to $n_{+}$, implying a conservative approach. Alternative, potentially less conservative treatments include dropping ties (adjusting $N$) or randomly breaking them. See also ? for another approach. The test has perfect power and does not produce false discoveries. Feature correlation (e.g. $\ell=1,6$) and discrete feature distributions ($\ell=9,10$) do not affect performance. The test also correctly detects the significance of features $\ell=2,4,5$ which influence the response only through product terms. Additional experiments show robustness to alternative baseline specifications (e.g. taking $v^{\ell}_{i}$ as the training mean also for the discrete features $\ell=9,10$, and using the median instead of the mean to construct $v^{\ell}_{i}$). We also verify robustness with respect to the score function $S$ including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
@@
We turn to measuring feature importance. Table 3 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen. It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities. Table 3 also provides the distribution of the randomized $p$-value. [^10] Ties of samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing CIs and coverage probabilities. This is consistent with their treatment in the testing procedure (footnote 8).
@@ 6.1 Credit card default
This section complements our theoretical and numerical results with an empirical application to credit card payment data from 30,000 credit card holders of a Taiwanese card issuer, collected in October 2005.131313The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients The response variable is a binary indicator of payment default. There are 23 explanatory variables (see Table 7). The dataset was split into training (60%), validation (15%), and testing (25%) sets with class proportions maintained through stratified sampling. Categorical variables are one-hot encoded and continuous variables are standardized (mean 0 and unit variance) based on training set statistics.
@@ 6.2 Mortgage risk
We use the feature effect grouping approach discussed in Section 2.2 for panel data, where a unit is a mortgage loan. The panel is unbalanced because loans enter the dataset at different times and are observed for different time periods. The negative multi-class cross-entropy loss is used as the score function $S$, leading to a test of feature importance jointly for transitions between all states. Alternative formulations are possible; for example, we could target specific transitions (e.g. current to prepayment). Baseline and comparison features are chosen as in Section 6.1, at the level of a group.
Table 10 summarizes the test results for $\alpha=1\%$ ($N=46,800$). In line with expectations, the most important feature by far is the current state of the mortgage. The market mortgage rate is the second most important feature, highlighting the significance of time-varying macro-economic conditions for mortgage borrower behavior. Indeed, borrowers have an incentive to prepay and refinance their mortgage if the market mortgage rate falls below their own current loan interest rate (the latter is ranked eighth). We also recognize path-dependency: the number of times the borrower was current and the number of times the borrower was 30 days delinquent during the past 12 months are highly significant and economically important variables for borrower behavior.
@@
rank variable median P(reject) lower upper 37.5% 62.5% coverage lower upper coverage
1 current_state 0.560 1.000 0.000 0.000 0.557 0.557 0.99 0.556 0.564 0.99016
2 national_mortgage_rate 0.006 1.000 5.372e-66 6.297e-66 0.005 0.005 0.99 0.005 0.006 0.99016
3 pool_insurance_flag_value 0.005 1.000 4.345e-314 6.182e-314 0.005 0.005 0.99 0.005 0.005 0.99016
4 times_current 0.002 1.000 5.920e-178 7.708e-178 0.002 0.002 0.99 0.002 0.003 0.99016
5 times_30dd_value 0.002 1.000 0.000 0.000 0.001 0.001 0.99 0.001 0.002 0.99029
6 loan_age 7.823e-04 1.000 3.830e-18 4.151e-18 5.401e-04 5.407e-04 0.99 5.230e-04 0.001 0.99016
7 lagged_prepayment_rate 7.518e-04 1.000 7.103e-127 8.870e-127 6.730e-04 6.732e-04 0.99 6.662e-04 8.298e-04 0.99029
8 current_interest_rate_value 6.497e-04 1.000 3.066e-83 3.667e-83 5.727e-04 5.730e-04 0.99 5.636e-04 7.351e-04 0.99016
9 inferred_collateral_type_value 5.229e-04 1.000 2.792e-43 3.173e-43 4.318e-04 4.318e-04 0.99 4.231e-04 6.206e-04 0.99016
10 mba_days_delinquent_value 3.738e-04 1.000 1.890e-131 2.370e-131 3.379e-04 3.380e-04 0.99 3.341e-04 4.136e-04 0.99016
11 initial_interest_rate 2.199e-04 1.000 1.219e-27 1.349e-27 1.775e-04 1.776e-04 0.99 1.751e-04 2.671e-04 0.99016
12 times_foreclosure 2.032e-04 1.000 2.492e-46 2.844e-46 1.680e-04 1.681e-04 0.99 1.631e-04 2.454e-04 0.99029
13 original_balance 1.815e-04 1.000 4.651e-43 5.283e-43 1.505e-04 1.505e-04 0.99 1.479e-04 2.178e-04 0.99016
14 prepay_penalty_flag_value 1.568e-04 1.000 6.926e-19 7.520e-19 1.152e-04 1.152e-04 0.99 1.115e-04 2.001e-04 0.99016
15 buydown_flag_value 1.294e-04 1.000 3.980e-35 4.462e-35 1.062e-04 1.064e-04 0.99 1.035e-04 1.536e-04 0.99016
16 fico_score_at_origination 1.233e-04 1.000 4.995e-36 5.609e-36 9.650e-05 9.669e-05 0.99 9.399e-05 1.515e-04 0.99016
17 scheduled_monthly_pi_value 3.479e-05 1.000 2.451e-30 2.726e-30 2.661e-05 2.663e-05 0.99 2.576e-05 4.341e-05 0.99016
- original_ltv 1.874e-05 0.000 0.017 0.018 -2.449e-06 -2.418e-06 0.99 -4.904e-06 4.245e-05 0.99016
- unemployment_rate -3.907e-04 0.000 1.000 1.000 -4.822e-04 -4.813e-04 0.99 -4.909e-04 -2.838e-04 0.99016
- current_balance_value -1.298e-06 0.000 0.520 0.524 -3.206e-05 -3.195e-05 0.99 -3.507e-05 3.639e-05 0.99016
- scheduled_principal_value -7.242e-06 0.000 1.000 1.000 -1.134e-05 -1.133e-05 0.99 -1.176e-05 -3.263e-06 0.99016
- convertible_flag_value -1.886e-04 0.000 1.000 1.000 -2.477e-04 -2.476e-04 0.99 -2.533e-04 -1.270e-04 0.99016
- io_flag_value -2.458e-04 0.000 1.000 1.000 -2.948e-04 -2.946e-04 0.99 -2.989e-04 -2.044e-04 0.99016
- negative_amortization_flag_value 9.342e-05 0.000 0.015 0.015 -7.297e-06 -6.803e-06 0.99 -1.626e-05 2.213e-04 0.99016
- original_term_value -1.047e-04 0.000 0.999 0.999 -1.854e-04 -1.852e-04 0.99 -1.936e-04 -1.765e-05 0.99016
- times_60dd -3.928e-04 0.000 1.000 1.000 -5.032e-04 -5.024e-04 0.99 -5.118e-04 -2.708e-04 0.99016
- times_90dd -4.781e-04 0.000 1.000 1.000 -5.630e-04 -5.630e-04 0.99 -5.723e-04 -3.823e-04 0.99016
- lagged_foreclosure_rate -1.049e-04 0.000 0.998 0.998 -1.769e-04 -1.765e-04 0.99 -1.846e-04 -9.468e-06 0.99016
Table 10: AICO test summary for the mortgage risk dataset with set-sequence model. Training data is now grouped and treated at sample-loan level. Continuous features are aggregated with mean while discrete and categorical features are aggregated by mode (just basic mode).
@@ 7 Conclusion
This paper develops a simple, statistically rigorous, and computationally efficient framework for assessing feature significance in supervised learning. Our randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation. Empirical studies show strong performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, demonstrating its ability to inform high-stakes decisions with legal, economic, and regulatory implications. More broadly, the framework lays the groundwork for a new standard of interpretability in machine learning—one that combines rigor with scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.

### v3

@@ 2.1 Baseline and comparison features
A few basic principles guide the selection of reference points and the construction of the set $\underline{I}\subseteq I_{1}$. We wish to obtain values $\underline{X}_{i}$ and $\overline{X}_{i}$ that are sufficiently distinct, to avoid feature effects $\Delta_{i}$ that are biased low which result in significance tests with low power.555? discuss these issues in the context of image classification. Similarly, we want to avoid saturation, where the predicted response is relatively flat in the neighborhood of $\underline{X}_{i}$. Finally, we might want to avoid the selection of values $(\underline{X}_{i},\overline{X}_{i})$ that are off the data manifold or in low-density areas of the training space. It is important to select an appropriate reference point and rule used to aggregate across $\underline{I}$ (see Example 2.1). Alternatively, we might select an appropriate baseline value from the set $\{X_{k}:k\in I_{1}\}$, namely an element that is near the raw off-manifold value emerging from the baseline selection procedure. [^6] We also allow for the selection and aggregation of alternative baselines for a given sample. A general tool is sample-dependent or feature-dependent selection of $(\underline{X}_{i},\overline{X}_{i})$.
2.2 Multiple features, panel data and missing data
We have focused on selecting $(\underline{X}_{i},\overline{X}_{i})$ for testing a single feature $\ell\in[d]$. The approach naturally extends to testing multiple features, i.e. subsets of $[d]$, which facilitates the exploration of higher-order effects including feature interactions.
Panel data settings, where a collection of units are observed over some time period, are prevalent in many domains. Appendix B discusses two simple extensions—conditioning and grouping—to construct the feature effect for panel data.
Real-world data sets often have missing feature values. Appendix C shows how to treat missing or imputed values when constructing $\Delta_{i}$.
@@ 5.1 Regression
We run 10 independent training/testing experiments to assess the behavior of the test. The second and third columns of Table 1 report the results for confidence levels $\alpha=0.01$ and 0.050.05.999Limited numerical precision can lead to ties, i.e. samples $\Delta_{i}$ whose true values are distinct are being rounded to the same value. If the value of a tie is strictly positive, we count each sample in the test statistic $n_{+}$. Another numerical artifact is samples $\Delta_{i}$ with true, non-zero value being rounded to 0. We treat these samples as if they indeed had non-positive values not contributing to $n_{+}$, implying a conservative approach. Alternative, potentially less conservative treatments include dropping ties (adjusting $N$) or randomly breaking them. See also ? for another approach. The test has perfect power and does not produce false discoveries. Feature correlation (e.g. $\ell=1,6$) and discrete feature distributions ($\ell=9,10$) do not affect performance. The test also correctly detects the significance of features $\ell=2,4,5$ which influence the response only through product terms. Additional experiments show robustness to alternative baseline specifications (e.g. taking $v^{\ell}_{i}$ as the training mean also for the discrete features $\ell=9,10$, and using the median instead of the mean to construct $v^{\ell}_{i}$). We also verify robustness with respect to the score function $S$ including the alternative specification S(m,y)=|my|S(m,y)=-|m-y|.
@@
We turn to measuring feature importance. Table 3 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen. It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities. Table 3 also provides the distribution of the randomized $p$-value. [^11] Ties of samples $\Delta_{i}$ due to limited numerical precision are retained in $\{\Delta_{i}:i\in I_{2}\}$ when computing CIs and coverage probabilities. This is consistent with their treatment in the testing procedure (footnote 9).
@@ 6.1 Credit card default
This section complements our theoretical and numerical results with an empirical application to credit card payment data from 30,000 credit card holders of a Taiwanese card issuer, collected in October 2005.141414The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients The response variable is a binary indicator of payment default. There are 23 explanatory variables (see Table 7). The dataset was split into training (60%), validation (15%), and testing (25%) sets with class proportions maintained through stratified sampling. Categorical variables are one-hot encoded and continuous variables are standardized (mean 0 and unit variance) based on training set statistics.
@@ 6.2 Mortgage risk
The dataset contains missing feature values, which are handled using missingness dummies and forward-filling based on the last observed value for each loan. To avoid biased test results, loan–month samples with missing values are excluded from the test set, as discussed in Appendix C. Because the extent of missingness varies across features, each feature test is conducted on a different effective sample size.
We apply the feature-effect grouping approach described in Appendix B for panel data, treating each mortgage loan as a unit. The panel is unbalanced because loans enter the dataset at different times and are observed for varying durations. The score function $S$ is the negative mean multi-class cross-entropy loss, where the mean is computed across time steps for each loan. This multi-period specification of $S$ enables a joint test of feature importance across all state transitions. Alternative formulations are also possible—for instance, targeting specific transitions such as from Current to Prepaid.
Baseline features are constructed through a two-step aggregation process: first from the loan–month level to the loan level, and then across all loans. For each feature $\ell$ and each loan $k\in I_{1}$ in the training set, the loan-level statistic $u_{k}^{\ell}$ is computed from the loan’s training records $\{X_{k,\cdot}:(k,\cdot)\in I_{1}\}$. Continuous variables are aggregated using the mean, whereas discrete and categorical variables use the mode. Using these loan-level aggregates, baseline and comparison features are then constructed according to (5.1), where $v_{i}^{\ell}$ is the mean of $\{u_{k}^{\ell}:k\in I_{1}\}$ for continuous features. For discrete and categorical variables, we instead use the alternative mode $v_{i}^{\ell}=\arg\max_{u\in\mathcal{X}^{\ell}\setminus\{u_{i}^{\ell}\}}p^{\ell}(u\mid I_{1})$, where $\mathcal{X}^{\ell}$ denotes the marginal support of $X^{\ell}$ and $p^{\ell}(\cdot\mid I_{1})$ is the empirical probability mass function of $\{u_{k}^{\ell}:k\in I_{1}\}$. This two-step aggregation ensures uniform weighting across loans when constructing baseline values, regardless of the number of observations available for each loan.
Table 10 summarizes the test results for $\alpha=1\%$. As expected, the most important feature by far is the current state of the mortgage. The market mortgage rate is the second most important feature, underscoring the influence of time-varying macroeconomic conditions on borrower behavior. Borrowers have a strong incentive to prepay or refinance their mortgage when the market rate falls below their own loan rate, which is ranked nineth in importance. The results also reveal pronounced path dependence: the number of months a borrower remained current and the number of months 30 days delinquent during the past 12 months are both highly significant and economically important predictors of borrower behavior.
@@
rank variable median P(reject) lower upper lower-1 lower-2 P(lower-2) coverage lower upper coverage missing rate
1 current_state 0.560 1.000 0.000 0.000 0.557 0.557 0.625 0.99 0.556 0.564 0.9902 0.000
2 national_mortgage_rate 0.006 1.000 5.372e-66 6.297e-66 0.005 0.005 0.625 0.99 0.005 0.006 0.9902 0.000
3 pool_insurance_flag_value 0.005 1.000 1.031e-240 1.560e-240 0.005 0.005 0.032 0.99 0.005 0.006 0.9903 0.443
4 times_current 0.002 1.000 5.920e-178 7.708e-178 0.002 0.002 0.625 0.99 0.002 0.003 0.9902 0.000
5 times_30dd_value 0.002 1.000 0.000 0.000 0.001 0.001 0.393 0.99 0.001 0.002 0.9902 0.014
6 loan_age 7.823e-04 1.000 3.830e-18 4.151e-18 5.401e-04 5.407e-04 0.625 0.99 5.230e-04 0.001 0.9902 0.000
7 inferred_collateral_type_value 7.683e-04 1.000 2.425e-70 2.882e-70 6.749e-04 6.750e-04 0.535 0.99 6.652e-04 8.658e-04 0.9902 0.096
8 lagged_prepayment_rate 7.518e-04 1.000 7.103e-127 8.870e-127 6.730e-04 6.732e-04 0.625 0.99 6.662e-04 8.298e-04 0.9903 0.000
9 current_interest_rate_value 6.571e-04 1.000 8.549e-85 1.024e-84 5.789e-04 5.792e-04 0.530 0.99 5.716e-04 7.453e-04 0.9902 0.003
10 mba_days_delinquent_value 3.738e-04 1.000 1.890e-131 2.370e-131 3.379e-04 3.380e-04 0.625 0.99 3.341e-04 4.136e-04 0.9902 0.000
11 prepay_penalty_flag_value 2.733e-04 1.000 4.209e-56 4.903e-56 2.365e-04 2.366e-04 0.078 0.99 2.317e-04 3.184e-04 0.9901 0.085
12 initial_interest_rate 2.199e-04 1.000 1.219e-27 1.349e-27 1.775e-04 1.776e-04 0.625 0.99 1.751e-04 2.671e-04 0.9902 0.000
13 buydown_flag_value 2.192e-04 1.000 1.214e-68 1.461e-68 1.882e-04 1.883e-04 0.160 0.99 1.856e-04 2.537e-04 0.9903 0.235
14 times_foreclosure 2.032e-04 1.000 2.492e-46 2.844e-46 1.680e-04 1.681e-04 0.625 0.99 1.631e-04 2.454e-04 0.9903 0.000
15 original_balance 1.815e-04 1.000 4.651e-43 5.283e-43 1.505e-04 1.505e-04 0.625 0.99 1.479e-04 2.178e-04 0.9902 0.000
16 fico_score_at_origination 1.233e-04 1.000 4.995e-36 5.609e-36 9.650e-05 9.669e-05 0.625 0.99 9.399e-05 1.515e-04 0.9902 0.000
17 scheduled_monthly_pi_value 3.616e-05 1.000 4.464e-32 4.979e-32 2.805e-05 2.809e-05 0.067 0.99 2.688e-05 4.472e-05 0.9900 0.003
- original_ltv 1.874e-05 0.000 0.017 0.018 -2.449e-06 -2.418e-06 0.625 0.99 -4.904e-06 4.245e-05 0.9902 0.000
- unemployment_rate -3.907e-04 0.000 1.000 1.000 -4.822e-04 -4.813e-04 0.625 0.99 -4.909e-04 -2.838e-04 0.9902 0.000
- current_balance_value -1.298e-06 0.000 0.520 0.524 -3.206e-05 -3.195e-05 0.625 0.99 -3.507e-05 3.639e-05 0.9902 0.000
- scheduled_principal_value -5.314e-05 0.000 1.000 1.000 -5.989e-05 -5.983e-05 0.980 0.99 -6.061e-05 -4.456e-05 0.9901 0.573
- convertible_flag_value -5.164e-04 0.000 1.000 1.000 -6.553e-04 -6.551e-04 0.034 0.99 -6.722e-04 -3.525e-04 0.9900 0.555
- io_flag_value -4.553e-05 0.000 0.973 0.973 -1.026e-04 -1.020e-04 0.868 0.99 -1.091e-04 1.431e-05 0.9902 0.324
- negative_amortization_flag_value -8.336e-04 0.000 1.000 1.000 -0.001 -0.001 0.825 0.99 -0.001 -4.979e-04 0.9903 0.853
- original_term_value -1.041e-04 0.000 0.999 0.999 -1.843e-04 -1.843e-04 0.139 0.99 -1.924e-04 -1.752e-05 0.9900 1.066e-04
- times_60dd -3.928e-04 0.000 1.000 1.000 -5.032e-04 -5.024e-04 0.625 0.99 -5.118e-04 -2.708e-04 0.9902 0.000
- times_90dd -4.781e-04 0.000 1.000 1.000 -5.630e-04 -5.630e-04 0.625 0.99 -5.723e-04 -3.823e-04 0.9902 0.000
- lagged_foreclosure_rate -1.049e-04 0.000 0.998 0.998 -1.769e-04 -1.765e-04 0.625 0.99 -1.846e-04 -9.468e-06 0.9902 0.000
Table 10: AICO test summary for the mortgage risk dataset using the set-sequence model. Training data is grouped and treated at the loan-month level. The missing rate denotes the ratio of feature effects dropped from the test set due to missigness to the total number of loan-level feature effects after grouping (46,890). P(lower-2) represents the probability associated with the larger lower bound (lower-2) of the one-sided randomized CI.
@@ 7 Conclusion
This paper develops a simple and computationally efficient framework for assessing feature significance in supervised learning. Our randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation. Empirical studies show strong performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, demonstrating its ability to inform high-stakes decisions with legal, economic, and regulatory implications. More broadly, the framework lays the groundwork for a new standard of interpretability in machine learning—one that combines statistical grounding with scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.